The vapour pressure of a sample of benzene is 53.0 kPa
at 60.6°C, but it fell to 51.2 kPa when 0.133 g of an organic
compound was dissolved in 5.00 g of the solvent. Calculate
the molar mass of the compound.
1
Expert's answer
2020-06-29T07:47:00-0400

Raoult's law states that the vapor pressure of a solvent above a solution is equal to the vapor pressure of the pure solvent at the same temperature scaled by the mole fraction of the solvent present:

Psolution = χ solvent Posolvent

χ solvent = Psolution / Posolvent = 51.2 / 53.0 = 0.966


χ solvent = moles benzene / total moles

M (C6H6) = 78.11 g / mol

n = 5.00 / 78.11 = 0.064 mol


M (A) = x g / mol (unknown)

n = 0.133 / x


Put all data together:


"0.966 = 0.064 \/ (0.064 + 0.133\/x)"

From the equation:

x = 58.7 (g/mol) - molar mass of A